OpenStudy (anonymous):

If y=Arcsin(3x/4), then dy/dx is

myininaya (myininaya):

i will let him do it the easy way lol

OpenStudy (anonymous):

\[\frac{d}{dx}\arcsin(x)=\frac{1}{\sqrt{1-x^2}}\] there i said it!

myininaya (myininaya):

\[\frac{d}{dx}\arcsin(u)=\frac{u'}{\sqrt{1-u^2}}\]

OpenStudy (anonymous):

so you only need to replace u by \[\frac{3x}{4}\] and \[u'\] by \[\frac{3}{4}\]

OpenStudy (anonymous):

So would I come out with [3/\sqrt{16-9x^2}\]